OpenStudy (anonymous):

Please check my work? Rectangle A has an area of 4 − x^2. Rectangle B has an area of x^2 + 2x − 8. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work. Rectangle A 4 - x^2 (x + 2)(x - 2) Rectangle B x^2 + 2x - 8 (x + 4)(x - 2) (x + 2)(x - 2)/(x + 4)(x - 2) x + 2/x + 4 The ratio of the area of Rectangle A to the area of Rectangle B is x+2/x+4. Is it correct?

OpenStudy (anonymous):

Incorrect

OpenStudy (campbell_st):

nope your factoring of rectangle A is incorrect. is the area is 4 - x^2 it factors to (2 - x)(2 + x)

OpenStudy (anonymous):

4 - x^2 = (2+x)(2-x), NOT (x-2)(x+2)

OpenStudy (anonymous):

Rectangle A 4 - x^2 (2 + x)(2 - x) Rectangle B x^2 + 2x - 8 (x + 4)(x - 2) (2 + x)(2 - x)/(x + 4)(x - 2) 2 + x/x + 4 The ratio of the area of Rectangle A to the area of Rectangle B is 2+x/x+4. is this right?

OpenStudy (anonymous):

You have factored 4 - x^2 incorrectly. It should be (2 -x)(2 + x)

OpenStudy (campbell_st):

nope the scaling factor is negative (2 - x) and (x - 2) are not the same -(x - 2) can be cancelled with (x -2) but its -1

OpenStudy (anonymous):

Sorry, you are right, I didnt see that you factored it correctly.

OpenStudy (anonymous):

It cancels with a -1

OpenStudy (anonymous):

Where do I get the -1 though?

OpenStudy (campbell_st):

so you are looking at \[\frac{-(x -2)(x+2)}{(x-2)(x+4)}\] now remove the common factor the initial area equation for rectangle A can be written as \[-(x^2 - 4) = -(x-2)(x+2)\]

OpenStudy (anonymous):

I'm sorry... I still don't get it...

OpenStudy (anonymous):

What I (and others) are saying: the expression (x-2)/(2-x) or (2-x)/(x-2) = -1

OpenStudy (anonymous):

2-x and x-2 are NOT identical, so you just cant cancel them out so fast; but they will cancel as the expression equals - 1.

OpenStudy (anonymous):

x+ 2 and 2+ x are identical, so they cancel out.

OpenStudy (anonymous):

I get that part I think, I just don't know how to write this all out

OpenStudy (campbell_st):

start from the basics write the ratio you need the rewrite the numerator so that you can get a common factor. \[\frac{4 - x^2}{x^2 + 2x - 8} = \frac{-(x^2 - 4)}{(x -2)(x + 4)}\]

OpenStudy (campbell_st):

which when factored is \[\frac{-1(x -2)(x+2)}{(x-2)(x+4)}\] only now can you cancel a common factor which leaves \[\frac{-(x +2)}{(x + 4)}\]

OpenStudy (anonymous):

Campbell is correct.

OpenStudy (anonymous):

Rectangle A 4 - x^2 -(x^2 - 4) -(x-2)(x + 2) Rectangle B x^2 + 2x - 8 (x + 4)(x - 2) -(x - 2)(x + 2)/(x + 4)(x - 2) x + 2/x + 4 The ratio of the area of Rectangle A to the area of Rectangle B is x+2/x+4. Is this one right?

OpenStudy (anonymous):

-(x+2)/(x+4)

OpenStudy (anonymous):

Kind of strange to have a negative ratio in the context of this problem.

OpenStudy (anonymous):

But thats what it is algebraically.

OpenStudy (anonymous):

Assuming that you typed the original question correctly.

OpenStudy (anonymous):

Rectangle A 4 - x^2 -(x^2 - 4) -(x-2)(x + 2) Rectangle B x^2 + 2x - 8 (x + 4)(x - 2) -(x - 2)(x + 2)/(x + 4)(x - 2) -(x + 2)/(x + 4) The ratio of the area of Rectangle A to the area of Rectangle B is -(x+2)/(x+4). How about this one??

OpenStudy (anonymous):

I agree with that.

OpenStudy (anonymous):

it's right??

OpenStudy (anonymous):

I said, that I would agree with that.

OpenStudy (anonymous):

Which means that you are right.

OpenStudy (anonymous):

haha okay! Well thank you for everything!!

OpenStudy (anonymous):

No problem.